An isosceles triangle is inscribed in a circle. Find the radius of the circle if one leg of the triangle is 8 cm. with 45 degree to 45 degree -90 degree. Please help me :)

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

An isosceles triangle is inscribed in a circle. Find the radius of the circle if one leg of the triangle is 8 cm. with 45 degree to 45 degree -90 degree.
===================
It's a right triangle, so the hypotenuse is the diameter.
Use Pythagoras to find the hypotenuse.

Answer by rothauserc(4718)

radius(r) = hypotenuse(h) of right isosceles triangle / 2
h^2 = 8^2 + 8^2
h^2 = 64 + 64 = 128
h = square root(128) = 11.31
r = 11.31 / 2 = 5.655
note. To see how this works imagine a square inscribed in the circle and draw one of the squares diagonals, this gives us two inscribed isosceles right triangles, we need to work with just one of them.